A college student sent a postcard to her parents with the message SEND + MORE = MONEY if each letter represents a digit (0, 1, 2, 3, 4, 5, 6, 7, 8, or 9), with different letters representing different digits and the same letter representing the same digits each time it occurs, how much money is being requested?


1
Expert's answer
2021-10-05T23:17:49-0400

"\frac{\begin{matrix}
&S&E&N&D\\
+
\\
&M&O&R&E\\
\end{matrix}}{\begin{matrix}
 M&O&N&E&Y
\end{matrix}}"


1) SEND+MORE<9999+9999=19998<20000. So, we must have M=1.

2) Now we have two cases: S+M+1=10+O or S+M=10+O.

S=8+O or S=9+O

O can’t be equal to 1 (we have M=1) and S"\leq" 9. Therefore, O=0 and S=8 or S=9.

If S=8, then E+O=10+N or E+O+1=10+N.

E=10+N or E=9+N

Since N"\geq" 2, we have that S"\neq" 8.

So, S=9.

3) E+O=N or E+O+1=N

E=N or E+1=N

Since E and N are different letters, we have E+1=N.

We can get it, if N+R=10+E or N+R+1=10+E.

So, R=9 or R=8. Since S=9, it follows that R=8.

4) We got R=8 from N+R+1=10+E. It means, that D+E=10+Y.

D+E"\leq" 6+7=13. So, 2"\leq" Y"\leq3" .

If Y=3, then (E=6 and D=7) or (E=7 and D=6).

If E=6, then N=E+1=7=D. If E=7, then N=E+1=8=R.

We have that Y=2.

5) If Y=2, then D+E=12.

3"\leq" D"\leq7"

D=3, E=9=S

D=4, E=8=R

D=5, E=7, N=E+1=8=R

D=6, E=6=D

D=7, E=5, N=E+1=6


M=1, O=0, S=9, R=8, Y=2, D=7, E=5, N=6


ANSWER: 9567+1085=10652, MONEY=10652.